# CBAM-EfficientNetV2 for Histopathology Image Classification using Transfer Learning and Dual Attention Mechanisms

### v4

@@ II Literature Review
Furthermore, Deformable Attention Mechanisms have recently gained traction for their adaptability in handling heterogeneous medical datasets. He et al. [36] introduced deformable attention layers that dynamically adjust receptive fields, outperforming traditional attention mechanisms in scenarios with high inter-patient variability, such as BreakHis and CAMELYON17 datasets.
@@ III-A Model Architecture
Our proposed model is based on EfficientNetV2-XL which use compound scaling to optimize accuracy and efficiency, making them ideal for handling the variability in histopathological images, such as differences in cellular structures and staining quality.
To enhance feature extraction, we integrate the Convolutional Block Attention Module (CBAM). CBAM combines channel and spatial attention, allowing the model to focus on informative features and key regions within the image. This is particularly beneficial in histopathology, where precise identification of cancerous structures is crucial. The EfficientNet backbone extracts initial features, and CBAM refines these representations before passing through fully connected layers for classification.
This integrated architecture, illustrated in Fig. 1, optimizes both efficiency and accuracy in breast cancer histopathology classification.
@@ III-B1 Channel Attention Module
1. Pooling Operations: The input feature map $F$ is first subjected to average pooling and maximum pooling to capture the global contextual information:
@@
Figure 2: Model Architecture
@@ III-E Training and Hyperparameter Tuning
The training of the models was conducted using PyTorch on CUDA, enabling efficient computation. We utilized an initial learning rate of 0.001 and employed Adam optimizers to determine which provided better convergence and generalization.
@@
  • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier.
  • Activation Functions:ReLU activation functions was used at the last classification layer.
  • Kernel Initializers: We experimented with different kernel initializers, including the normal initializer and He initialization, to determine their effects on the convergence rate during training.
  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4.
@@ III-F Evaluation Metrics
Figure 5: Visualization of Attention Mechanisms in Histopathological Images
The experiments were conducted using PyTorch as the deep learning framework backend. The models were trained and tested on an NVIDIA RTX 4080 GPU, providing substantial computational power for the deep learning tasks involved. The system was equipped with 16GB of RAM to efficiently handle the large dataset and facilitate faster processing.
@@
We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate ( $0.001$. The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
Binary cross-entropy was used as the loss function due to the binary nature of our classification problem, represented mathematically as follows:
@@ IV-A Results
Initially, we conducted experiments on two different models without transfer learning, as presented in Table 3. Subsequently, we evaluated state-of-the-art models pre-trained on ImageNet, as summarized. The performance of our proposed modality-specific strategy, incorporating transfer learning.
As shown in Table 2, our proposed EfficientNetV2-XL + CBAM significantly outperformed various established models, including DenseNet201 and Deep CNNs. Notably, at higher magnification levels, this model achieved an outstanding accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, highlighting its ability to accurately classify breast cancer histopathology images
The EfficientNetV2-XL + CBAM demonstrated substantial performance improvements over the VGG16 + VGG19 combination across all magnifications. While the VGG-based models delivered competitive accuracy, their computational requirements were significantly higher. For instance, the VGG16 + VGG19 models required a Tesla K80 GPU with 24GB of VRAM to accommodate their large parameter count and high computational demand.
In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4080 GPU with just 16GB of VRAM. Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
@@ V Conclusion
In this study, we propose a Hybrid EfficientNetV2 models integrated with CBAM, Self-Attention, and Deformable Attention mechanisms for breast cancer histopathology classification. The EfficientNetV2-XL model, with approximately 208 million parameters and the inclusion of the CBAM module, offers a more efficient and accurate solution compared to previous state-of-the-art models like VGG16 + VGG19, which have around 282 million parameters. This reduction in parameters enables faster training and reduced memory consumption, allowing the model to be trained on an RTX 4060 GPU with 8GB of memory.
The Hybrid EfficientNetV2-XL + CBAM and previous State-of-the-Art model VGG16 + VGG19 model achieved superior accuracy, however EfficientNetv2-XL has F1-Score of 98.31 and Hybrid VGG16 + VG199 has 97.95% ,therefore EfficientNetV2-XL + CBAM outperforms other models in both accuracy and computational efficiency. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment. Future work may adapt the model to other datasets and further expand its clinical applicability.

### v5

@@ III-A Model Architecture
EfficientNetV2-XL employs a compound scaling strategy to optimize the network’s depth, width, and resolution, making it particularly well-suited for medical image classification tasks where diverse cellular structures and magnification levels exist. This backbone allows the network to adjust the number of convolutional blocks based on the complexity of the input image. Visualization of the Architecture in Figure 1.
  • Conv Block 1:
    • 3x3 convolutions with stride 1 and padding to capture spatial features at a finer resolution, followed by ReLU activation to introduce non-linearity and enhance feature learning.
    • Followed by 1x1 convolutions to reduce channel dimensions while preserving essential spatial features, also followed by ReLU activation.
    • Batch Normalization (BN) is applied after each convolution to stabilize training and improve generalization.
    • The block concludes with 2x2 max pooling, which reduces the spatial dimensions while preserving key features.
  • Conv Block 2:
    • Consolidates further layers from the EfficientNet backbone with additional 3x3 convolutions for deep feature extraction, each followed by ReLU activation.
    • Followed by 1x1 convolutions to maintain a computationally efficient structure, paired with ReLU activation.
    • Max pooling is applied again to downsample the feature maps, allowing for multi-scale feature representation.
  • Conv Block 3:
    • The final block retains a 3x3 convolution with ReLU activation but without further pooling to maintain spatial resolution at this stage.
    • BN is also applied here, followed by Global Average Pooling (GAP), which reduces the feature maps to a single value per feature map while retaining global information.
Each block’s output is passed through a CBAM module, which refines the feature maps by applying channel and spatial attention mechanisms, helping the model focus on the most critical parts of the input image.
@@ III-B1 Channel Attention Module
1. Pooling Operations: The input feature map $F$ is first subjected to average and maximum grouping to capture global contextual information:
@@
Figure 3: Spatial and Channel Attention
@@ III-E Training and Hyperparameter Tuning
We employed the Adam optimizer with a cosine annealing learning rate schedule and warm restarts. The initial learning rate ($\eta_{\text{max}}$) was set to $0.001$, with a minimum learning rate ($\eta_{\text{min}}$) of $0.0001$. The warm restart period ($T_{0}$) was initialized at 10 epochs, doubling with each subsequent restart (e.g., $T_{0}=10$, $T_{1}=20$, $T_{2}=40$). The optimizer parameters were tuned as follows: $\beta_{1}=0.9$, $\beta_{2}=0.98$, and $\epsilon=10^{-8}$. The learning rate at epoch $t$ was calculated using:
$\eta_{t}=\eta_{\text{min}}+\frac{1}{2}\left(\eta_{\text{max}}-\eta_{\text{min}}\right)\left(1+\cos\left(\frac{t}{T}\pi\right)\right)$
@@
  • Dense Layers: Experiments were conducted with two and three dense layers to identify the optimal configuration for the final classifier and settled upon two layers.
  • Activation Functions: SoftMax activation function was used at the last classification layer.
  • Kernel Initializers: We experimented with different kernel initializers, and settled with He initialization,
  • Dropout Regularization: Dropout was incorporated in the fully connected layers to prevent overfitting, particularly given the limited availability of labeled data in medical imaging tasks with a rate of 0.4 after experimenting with different values.
@@ III-F Evaluation Metrics
The experiments were conducted using PyTorch as the deep learning framework backend. The models were trained and tested on an NVIDIA RTX 4080 GPU, providing substantial computational power for the deep learning tasks involved. The system was equipped with 16GB of VRAM to efficiently handle the large dataset and facilitate faster processing.
@@
We evaluated two different optimizers: Adam (Adaptive Moment Estimation), with learning rate 0.001. The SoftMax classifier was employed for all classification experiments, providing probability scores for each class label.
Categorical Cross-entropy was used as the loss function due to the nature of our classification problem, represented mathematically as follows:
@@ IV-A Results
To further validate the effectiveness of our proposed CBAM-EfficientNetV2 model, we conducted qualitative analyses using attention maps. These visualizations illustrate the regions the model focuses on during classification, providing insight into its decision-making process.
The CBAM module enhances the model’s ability to highlight critical regions, such as irregular nuclei shapes and abnormal cellular arrangements, which are key indicators of malignancy.
The EfficientNetV2-XL + CBAM demonstrated substantial performance improvements over the VGG16 + VGG19 combination across all magnifications. While the VGG-based models delivered competitive accuracy, their computational requirements were significantly higher. For instance, the VGG16 + VGG19 models required a Tesla K80 GPU with 24GB of VRAM to accommodate their large parameter count and high computational demand. In contrast, the EfficientNetV2-XL + CBAM was trained on a more modest NVIDIA RTX 4080 GPU with just 16GB of VRAM. Despite the lower memory availability, the EfficientNetV2-XL architecture demonstrated exceptional computational efficiency, leveraging depthwise separable convolutions, squeeze-and-excitation (SE) blocks, and the Convolutional Block Attention Module (CBAM).
@@ V Conclusion
CBAM-EfficientNetV2-XL achieved superior accuracy of 99.01%, however EfficientNetv2-XL has F1-Score of 98.31 ,therefore EfficientNetV2-XL + CBAM outperforms other models in accuracy. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment. Future work may adapt the model to other datasets and further expand its clinical applicability.